how many quarts of pure antifreeze must be added to 5 quarts of a 10% antifreeze solution to obtain a 50% antifreeze solution?
Answer by richwmiller(17219)

.1*5+x=.5(5+x)
.5+x=2.5+.5x
.5x=2
x=4
check
.5+4=2.5+2
4.5=4.5
ok